OpenStudy (anonymous):

Langston will pay $575 to three different credit cards: Expert Card, Passport, and US Speedy. She will pay three times as much to Expert Card as US Speedy since it has a higher balance. She also expects the Passport payment to be only 15% of the total of the other two payments. How much will Langston pay towards the Passport credit card balance? $75.00 $86.25 $125.00 $375.00

OpenStudy (mathstudent55):

Let the amount paid to Expaert = e Let the amount paid to Passport = p Let the amount paid to US Speedy = u e + p + u = 575 e = 3u p = 0.15(e + u), but e = 3u, so p = 0.15(3u + u) = 0.15(4u) = 0.6u Substitute 3u for e and 0.6u for p and solve for u. Then p = 0.6u

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

What variables go into u or p or e???

OpenStudy (mathstudent55):

u, p, and e are the variables. I'll explain again more slowly.

OpenStudy (mathstudent55):

The total amount paid out is $575. Out of that, part will go to Expert, part to Passport, and part to US Speedy. Ok?

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

I am using the variables e, p, and u, to represent the amounts paid to Expert, Passport and US Speedy, respectively. Ok?

OpenStudy (mathstudent55):

The total paid to all three cards is $575. That means e + p + u = 575. Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now we need the info they give us about how much is paid to each card.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

She will pay three times as much to Expert Card as US Speedy From here we get e = 3u ok?

OpenStudy (anonymous):

then?

OpenStudy (mathstudent55):

From here: "She also expects the Passport payment to be only 15% of the total of the other two payments." we get p = 15%(e + u) ok?

OpenStudy (anonymous):

go on

OpenStudy (mathstudent55):

This expression for p can be simplified. p = 15%(e + u) We know that e = 3u, so let's replace e with 3u: p = 15%(3u + u)

OpenStudy (mathstudent55):

This can be simplified further: p = 15%(3u + u) p = 15%(4u) p = 0.15(4u) p = 0.6u

OpenStudy (anonymous):

Maybe it'd be easier for me to understand if you actually did the problem without putting the answer so I can make sense of it myself, and if I have any questions I can ask.

OpenStudy (mathstudent55):

We didn't get to the answer yet. We are just getting there slowly.

OpenStudy (mathstudent55):

Now we go back to e + p + u = 575 Since e = 3u, and p = 0.6u, we susbtitute in the equation: e + p + u = 575 3u + 0.6u + u = 575 Now you can solve for u.

OpenStudy (anonymous):

and how do i solve for u?

OpenStudy (anonymous):

Also how did you get 0.6u

OpenStudy (mathstudent55):

Add 3u plus 0.6u plus u

OpenStudy (anonymous):

.....

OpenStudy (mathstudent55):

Read above my answer that starts "This can be simplified further:" for an explanation of 0.6u

OpenStudy (anonymous):

Would it be 3.6u^2 = 575 ?????????

OpenStudy (mathstudent55):

3u + 0.6u + u = 575 u is the same as 1u 3u + 0.6u + 1u = 575 4.6u = 575

OpenStudy (anonymous):

So u = 125

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

now e = 3u, so e = 3(125)

OpenStudy (mathstudent55):

p = 0.6u, so p = 0.6(125)

OpenStudy (anonymous):

So the answer would be $75

OpenStudy (mathstudent55):

Yes

OpenStudy (anonymous):

thank you!